OpenStudy (anonymous):

What is the simplified for of the expression SqrT(16x^3)-SqrT(c^2)+5SqrT(c^2)+SqrT(4c^2) 8c4 + 8c 4c4 + 8c 8c4 + 6c 4c4 + 6c

OpenStudy (anonymous):

sqrt(4x^2) break down the sqrt ^ 2 2 2 sqrt(x^2) is the answer

OpenStudy (anonymous):

do you understand :D

OpenStudy (anonymous):

No because for 1, that doesn't make sence to me. 2, thats not one of my answer choices.

OpenStudy (anonymous):

@callie2240

OpenStudy (anonymous):

oops srorry i ddnt see the rest of the problem give me a sec

OpenStudy (jhannybean):

\[\large \sqrt{16x^3}-\sqrt{c^2}+5\sqrt{c^2}+\sqrt{4c^2}\]\[\large \sqrt{4*4*x*x*x}-\sqrt{c^2}+5\sqrt{c^2}+\sqrt{4c^2}\]\[\large 4x^2\sqrt{x}-\sqrt{c^2}+5\sqrt{c^2}+2\sqrt{c^2}\]\[\large 4x^2\sqrt{x}-6\sqrt{c^2}\]\[\large 4x^2\sqrt{x}-6c\]

OpenStudy (jhannybean):

I think you have your variables mixed up because your answers have all c's and your problem has an x.

OpenStudy (anonymous):

SqrT(16x^3)-SqrT(c^2)+5SqrT(c^2)+SqrT(4c^2) sqrt(15x^3)-55sqrt(6c^2) is your answer

OpenStudy (anonymous):

wait no now break it down lol

OpenStudy (jhannybean):

@Loser66 Did i do this correctly? lol

OpenStudy (loser66):

hehehe why me?

OpenStudy (anonymous):

@Jhannybean habbit of putting x's . but your answer isnt one of my answer choices?

OpenStudy (loser66):

@Jhannybean +6c not -

OpenStudy (loser66):

and the first term is 4xsqrt(x) hehe.

OpenStudy (jhannybean):

Oh yes. Whoops. \[\large 4x \sqrt{x}+6c\] *** corrected my bottom portion,

OpenStudy (loser66):

better

OpenStudy (jhannybean):

\[\large \sqrt{16c^3}-\sqrt{c^2}+5\sqrt{c^2}+4\sqrt{c^2}\]\[\large 4c \sqrt{c}-\sqrt{c^2}+5\sqrt{c^2}+2\sqrt{c^2}\]\[\large 4c \sqrt{c}+6 \sqrt{c^2}\]\[\large 4c \sqrt{c}+6c\]

OpenStudy (loser66):

@Caitlynn_Faucher your problem is not right, first, at the first term, it should be c, not x. if it's x , you must have something else to link between x and c to get those kind of answer second, base on the problem, we solve it, separately,

OpenStudy (loser66):

as long as your logic is strong, no need to care about the choices.

OpenStudy (loser66):

the last post of Jhannybean is perfect.

OpenStudy (loser66):

@Jhannybean please, don't point me out, let this older be in his laziness.

OpenStudy (jhannybean):

\[\large 4c \sqrt{c}+6c\]\[\large 4c^{3/2}+6c\]\[\large 2c(2\sqrt{c}+3)\] factoring it out even further.

OpenStudy (loser66):

ok

OpenStudy (jhannybean):

But it's STILL not one of the answer choices, the factor would be ((2 rad c) +3) and the GFC would be 2c.

OpenStudy (loser66):

@Jhannybean read my comments

OpenStudy (loser66):

don't forget, every one has mistake, even computer.!!! sometimes has mistake, too